# disposition

### v1

@@ Abstract
Keywords: Vector optimization; efficiency; weak efficiency; proper efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
@@ 1 Introduction
The main Section 5 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces. The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions. In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector optimization problem (based on the convex cone $C$), an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions). Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorem 5.25 for proper efficiency w.r.t. a cone-valued map, Theorems 5.38, 5.40 for Henig proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
@@ 2.2 Cones and their bases
$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\quad\text{and}\quad(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
@@
Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by
$C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}.$
It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]). Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone. If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and
$C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset.$
Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$. It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 4.11). Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 2.3), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]).
@@
Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define a set
@@ 2.3 Augmented dual cones and Bishop-Phelps type cones
Consider some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. Then, we have:
@@
    Nontrivialness: If $(x^{*},\alpha)\in K^{a\#}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
@@
$3^{\circ}$ This assertion is a direct consequence of Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
@@
  • Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. According to [37, Ex. 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has
    $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$
    and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
@@
    Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, the weak interior of $C$ is empty, i.e., $\operatorname{int}_{w}\,C=\emptyset$.
@@
    On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$. Take some $c\in\operatorname{int}_{w}\,C$, i.e., there is a weakly open neighbourhood $U(c)\subseteq Y$ around $c$ which satisfies $U(c)\subseteq C$. Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$. This implies, for any $n\in\mathbb{N}$,
    $\frac{c}{n}\pm v=\frac{1}{n}(c\pm nv)\in C.$
    Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
@@ 2.4 Solution concepts in vector optimization
where we assume throughout the paper that $K\subseteq Y$ is a cone. It is well-known that $K$ induces on $Y$ a reflexive binary relation $\leq_{K}$ defined, for any two points $y,\overline{y}\in Y$, by $y\leq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K.$ Note that if $K$ is convex, then $\leq_{K}$ is a preorder; if $K$ is pointed and convex, then $\leq_{K}$ is a partial order. For notational convenience, we consider further the binary relations $\lneq_{K}$ and $<_{K}$ that are defined, for any two points $y,\overline{y}\in Y$, by $y\lneq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K\setminus\{0\},y<_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-\operatorname{int}\,K$, respectively. Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., Boţ, Grad and Wanka [5], Göpfert et al. [23], Jahn [42], and Khan, Tammer and Zălinescu [49]).
@@
A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$. The set of all Pareto efficient solutions of ($P$) is denoted by
${\mathrm{E}ff}(\Omega\mid f,K):=\{\overline{x}\in\Omega\mid\not\exists\,x\in\Omega:\;f(x)\lneq_{K}f(\overline{x})\}.$
@@
A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$. The set of all weakly Pareto efficient solutions of ($P$) is denoted by
${\mathrm{W}Eff}(\Omega\mid f,K):=\{\overline{x}\in\Omega\mid\not\exists\,x\in\Omega:\;f(x)<_{K}f(\overline{x})\}.$
@@
Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$. A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and
$\mathbb{A}(\bar{x})\cap(-K)=\{0\}.$
The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
@@ 3 Scalarization in Vector Optimization
where $\varphi:Y\to\mathbb{R}$ is a (frequently parameter-dependent) function. Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$. By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$). Applying such a strategy is called scalarization method in the literature of vector optimization. Within such methods, the function $\varphi$ is called scalarizing function. For more details, we also refer the reader to the standard books of vector and set optimization by Boţ, Grad and Wanka [5], Göpfert et al. [23], Jahn [42], and Khan, Tammer and Zălinescu [49].
@@
The following scalarization results involving cone-monotone functions are well-known.
@@
Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, the following assertions hold:
@@
Suppose that $K\subseteq Y$ is a convex cone. Then, the following assertions are valid:
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the scalar optimization problem (3.1).
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
@@
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($3^{\circ}$).
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of the scalar optimization problem (3.1).
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that
    $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
@@
$4^{\circ}$. The first equality is obvious, while the second equality follows immediately by $1^{\circ}$.

### v2

@@ Abstract
Keywords: Vector optimization; (proper, weak) efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
@@ 1 Introduction
The main Section 5 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces. The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions. In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions). Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorems 5.25, 5.38 and 5.40 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
@@ 2.2 Cones and their bases
$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
@@
Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by $C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}$. It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]). Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone. If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and $C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$. It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 4.11). Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 2.3), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]).
@@
Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define
@@ 2.3 Augmented dual cones and Bishop-Phelps type cones
Consider $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. The following assertions hold:
@@
    Nontrivialness: If $(x^{*},\alpha)\in K^{a+}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
@@
$3^{\circ}$ This assertion follows from Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
@@
Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. According to [37, Ex. 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$ and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
@@
Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, we have $\operatorname{int}_{w}\,C=\emptyset$.
@@
On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$. Take some $c\in\operatorname{int}_{w}\,C$, i.e., there is a weakly open neighbourhood $U(c)\subseteq Y$ around $c$ which satisfies $U(c)\subseteq C$. Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$. This implies, for any $n\in\mathbb{N}$, we have $cn^{-1}\pm v=n^{-1}(c\pm nv)\in C$. Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
@@ 2.4 Solution concepts in vector optimization
where we assume throughout the paper that $K\subseteq Y$ is a cone. It is well-known that $K$ induces on $Y$ a reflexive binary relation $\leq_{K}$ defined, for any two points $y,\overline{y}\in Y$, by $y\leq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K.$ Note that if $K$ is convex, then $\leq_{K}$ is a preorder; if $K$ is pointed and convex, then $\leq_{K}$ is a partial order. For notational convenience, we consider further the binary relations $\lneq_{K}$ and $<_{K}$ that are defined, for any two points $y,\overline{y}\in Y$, by $y\lneq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K\setminus\{0\},y<_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-\operatorname{int}\,K$, respectively. Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., [5], [23], [42], [49]).
@@
A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$. The set of all Pareto efficient solutions of ($P$) is denoted by ${\mathrm{E}ff}(\Omega\mid f,K)$.
@@
A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$. The set of all weakly Pareto efficient solutions of ($P$) is denoted by ${\mathrm{W}Eff}(\Omega\mid f,K)$.
@@
Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$. A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$. The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
@@ 3 Scalarization in Vector Optimization
where $\varphi:Y\to\mathbb{R}$ is a (frequently parameter-dependent) function. Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$. By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$). The application of such a strategy is called scalarization method, while the function $\varphi$ is called scalarizing function (see e.g. [5], [23], [42], [49]).
@@
The next scalarization results involving cone-monotone functions are well-known.
@@
Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, we have:
@@
Suppose that $K\subseteq Y$ is a convex cone. Then, we have:
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (3.1).
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
@@
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 3.8 ($3^{\circ}$).
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of (3.1).
@@
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
@@
$4^{\circ}$. The first equality is obvious, while the second equality follows by $1^{\circ}$.